# Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Knowledge Distillation and Random Data Erasing

### v2

@@ 3.2.1 Main Results
As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset. Specifically, it outperforms MulT by 0.92%, the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5. Similar trends are observed in other settings. In contrary, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8). This highlights the consistent supremacy of our NECHO v2.
This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation, MAG distillation, and dual-level logit distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by single point data random erasing to minimise data distribution gaps. These enables the student to imitate the teacher in varied incompleteness settings, ensuring considerable performance gain effectively.
@@ 3.2.2 Ablation Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 KD w/o $\mathcal{L}_{\text{MWCD}}$ 37.10 54.10 34.37 50.79 w/o $\mathcal{L}_{\text{TR2D}}$ 36.9 53.95 33.15 50.58 w/o $\mathcal{L}_{\text{MAGD}}$ 36.01 53.27 32.91 49.87 w/o $\mathcal{L}_{\text{hrchyLD}}$ 35.58 52.85 34.25 50.96 DA only distillation 36.42 53.32 34.05 50.83 only teacher training 37.28 53.65 32.71 49.72 not at all 36.43 53.68 33.54 50.93 NECHO v2 Full 37.02 54.26 34.24 51.01
@@
To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3. We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representation is either highly dominant or minimal.
We first assess the effectiveness of KD. Whilst $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance. The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision. Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial in performance. The quantitative results validate the importance of all components in systematic KD to align student’s semantic knowledge to that of the teacher.
We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all. Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information. This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in a student model that is highly resilient to uncertain data incompleteness.
@@ 3.2.3 Comparative Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 Pairing Original $\to$ Original 36.79 53.11 32.68 49.53 Original $\to$ Modified w/ MAG 36.44 52.96 34.66 50.99 $\mathcal{L}_{\text{TR2D}}$ Not Random 36.91 53.56 32.83 50.40 NECHO v2 Proposed 37.02 54.26 34.24 51.01
@@
As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified with MAG) and transformer full distillation under the same settings with the ablation studies. Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
@@ 4 Conclusion
We tackle uncertain missing sequences for robust multimodal SDP with the proposed NECHO v2. Extensive experiments show the stable superiority of NECHO v2 over the existing methods. The code will be released soon.

### v3

@@ 3.2.1 Main Results
As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missingness scenarios on MIMIC-III dataset. Specifically, it outperforms MulT by 0.92%, the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5. Similar trends are observed in other settings.
In contrast, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8). UnimodalKD and Missmodal underperform in most incomplete scenarios, highlighting the need for systematic knowledge distillation that accounts for fluctuating modality dominance under imperfect data.
This remarkable performance gain of NECHO v2 is attributed to: 1) modifying NECHO to manage varying modality significance under imperfect data, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by curriculum random data erasing to minimise data distribution gaps. These enables the student to imitate the teacher in varied incompleteness settings, ensuring considerable performance gain effectively.
@@ 3.2.2 Ablation Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 KD w/o $\mathcal{L}_{\text{MWCD}}$ 37.10 54.10 34.37 50.79 w/o $\mathcal{L}_{\text{TR2D}}$ 36.9 53.95 33.15 50.58 w/o $\mathcal{L}_{\text{MAGD}}$ 36.01 53.27 32.91 49.87 w/o $\mathcal{L}_{\text{hrchyLD}}$ 35.58 52.85 34.25 50.96 DA Only During Distillation 36.42 53.32 34.05 50.83 Only During Teacher Training 37.28 53.65 32.71 49.72 Not For Both 36.43 53.68 33.54 50.93 NECHO v2 Full 37.02 54.26 34.24 51.01
@@
To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3. We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representations are either highly dominant or minimal.
We first assess the effectiveness of KD. Whilst NECHO v2 occasionally performs better without $\mathcal{L}_{\text{MWCD}}$ and $\mathcal{L}_{\text{MWHD}}$, their consistent use generally enhances performance. The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate representation propagation. Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial. These validate the importance of all components in our systematic KD pipeline to align the student’s semantic knowledge to that of the teacher.
We also evaluate the efficacy of data erasing against three scenarios: only during distillation, only during teacher training, and not for both. Overall performance considerably improves, highlighting the significance of the proposed curriculum random data erasing under missing visit information. This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in the student model that is highly resilient to uncertain data incompleteness.
@@ 3.2.3 Comparative Studies
Criteria Components (0.2, 0.2, 0.2) (0.5, 0.2, 0.8) top-10 top-20 top-10 top-20 Pairing Original $\to$ Original 36.79 53.11 32.68 49.53 Original $\to$ Modified for Student 36.44 52.96 34.66 50.99 $\mathcal{L}_{\text{TR2D}}$ Not Random 36.91 53.56 32.83 50.40 NECHO v2 Proposed 37.02 54.26 34.24 51.01
@@
Under the same settings as the ablation studies, we compare our NECHO v2 with different teacher-student combinations (original to original, original to modified for student) and transformer not random distillation. Our proposed methodologies achieve the best overall performance, underscoring the importance of: 1) carefully pairing teacher and student to address shifting representation dominance and minimise architectural heterogeneity, and 2) incorporating randomness into KD to prevent overfitting. We provide the corresponding result to Table 4.
@@ 4 Conclusion
We tackle uncertain missing sequences for robust multimodal SDP with the proposed NECHO v2. With modified NECHO that dynamically adjusts dominant representations under varying missingness, we design a curriculum data erasing guided systematic KD pipeline that enables the student to effectively imitate the teacher. Extensive experiments on MIMIC-III data show the effectiveness of our approach over the existing methodologies. To foster future research, we release code at: https://www.github.com/heejkoo9/NECHOv2.